# Elicitability and identifiability of tail risk measures

### v2

@@ Abstract
Tail risk measures are fully determined by the distribution of the underlying loss beyond its quantile at a certain level, with Value-at-Risk and Expected Shortfall being prime examples. They are induced by law-based risk measures, called their generators, evaluated on the tail distribution. This paper establishes joint identifiability and elicitability results of tail risk measures together with the corresponding quantile, provided that their generators are identifiable and elicitable, respectively. As an example, we establish the joint identifiability and elicitability of the tail expectile together with the quantile. The corresponding consistent scores constitute a novel class of weighted scores, nesting the known class of scores of Fissler and Ziegel for the Expected Shortfall together with the quantile. For statistical purposes, our results pave the way to easier model fitting for tail risk measures via regression and the generalized method of moments, but also model comparison and model validation in terms of established backtesting procedures.
@@ 1 Introduction
In view of the practical importance of tail risk measures, this paper studies their identifiability and elicitability. A risk measure is identifiable if it is the unique zero (root) of an expected identification function. Similarly, it is elicitable if it is the unique minimizer of an expected score; see Definitions 3.1 and 3.2. As such, identifiability and elicitability facilitate Z- and M-estimation (Huber and Ronchetti, 2009). More generally, model fitting can be pursued via the generalized method of moments (Newey and McFadden, 1994) or regression (Dimitriadis et al., 2024), also using recent tools from machine learning; see Cont et al. (2022) and Fissler et al. (2023) for examples. For model comparison and model validation, commonly subsumed under the umbrella term backtesting, identifiability and elicitability are crucial (Nolde and Ziegel, 2017; Fissler et al., 2016).
@@ 2.1 Notation
We work with an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$. Let $\mathcal{M}^{0}$ be the set of all Borel probability distributions on $\mathbb{R}$ and let $\mathcal{M}^{q}\subseteq\mathcal{M}^{0}$, $q\in[1,\infty)$, be the subset of all distributions with a finite $q$th moment. Moreover, $\mathcal{M}^{\infty}$ denotes the set of all compactly supported distributions. For the ease of presentation, we identify distributions with the corresponding cumulative distribution functions; that is, $F(x)$ is understood as $F((-\infty,x])$ for $x\in\mathbb{R}$. For $F\in\mathcal{M}^{0}$, we define the left-continuous generalized inverse (left-quantile) as
@@
and in addition set $F^{-1}(0)=\inf\{x\in\mathbb{R}:F(x)>0\}$. For $x\in\mathbb{R}$, denote by $\delta_{x}$ the point-mass probability measure at $x$. For simplicity, $\int$ means integration on $\mathbb{R}$ unless otherwise specified. Throughout this article, if we do not specify $\mathcal{M}$, then the statements hold for any set $\mathcal{M}\subseteq\mathcal{M}^{0}$ such that the functional at consideration is well-defined and finite on $\mathcal{M}$.
@@ 2.2 Risk measures
A risk measure is a mapping $\hat{\rho}$ from a set of random variables $\mathcal{X}$ on $(\Omega,\mathcal{F},\mathbb{P})$ to $\mathbb{R}$ (Föllmer and Schied, 2016). The typical interpretation of a risk measure is that its argument, $X\in\mathcal{X}$, represents the loss of some financial position, and that $\hat{\rho}(X)\in\mathbb{R}$ represents the amount of capital required to hold to make the risk of $X$ acceptable (we follow the sign convention of McNeil et al. (2015)). In Appendix A, we list several well studied properties for risk measures on $\mathcal{X}$. Most common risk measures are law-based (also called law-invariant), meaning that for any $X,X^{\prime}\in\mathcal{X}$ it holds that $\hat{\rho}(X)=\hat{\rho}(X^{\prime})$ whenever the distributions of $X$ and $X^{\prime}$ coincide, i.e., $F_{X}=F_{X^{\prime}}$. Let =$\mathcal{M}=\{F_{X}\in\mathcal{M}^{0}:X\in\mathcal{X}\}$. Then, a law-based risk measure $\hat{\rho}:\mathcal{X}\to\mathbb{R}$ induces a mapping
$\rho\colon\mathcal{M}\to\mathbb{R},\qquad\rho(F):=\hat{\rho}(X)\quad\text{for some $X\in\mathcal{X}$ with $F_{X}=F$.}$ (2.1)
In this paper, we shall take the mapping $\rho$ as the primary object and also call it a risk measure, slightly overloading terminology. Clearly, invoking the relation (2.1), our study can easily be translated into properties of $\hat{\rho}$.
The two most popular classes of risk measures used in banking and insurance practice are the Value-at-Risk (VaR) and the Expected Shortfall (ES). For a confidence level $p\in(0,1)$, the right-quantile ($\mathrm{VaR}^{+}_{p}$) and the left-quantile ($\mathrm{VaR}_{p}^{-}$), defined for $F\in\mathcal{M}^{0}$ as
$\mathrm{VaR}^{+}_{p}(F)$ $=\inf\{x\in\mathbb{R}:F(x)>p\}=F^{-1}(p+),$
$\mathrm{VaR}^{-}_{p}(F)$ $=\inf\{x\in\mathbb{R}:F(x)\geqslant p\}=F^{-1}(p).$
@@ 2.3 Tail distribution and tail risk measures
Invoking our sign convention that a distribution $F\in\mathcal{M}^{0}$ is a loss distribution, it is clear that the this right tail is the region of interest from a risk measurement and management perspective (McNeil et al., 2015). Clearly, with another sign convention and from a mathematical perspective, we could similarly consider the left tail of the distribution, and all the results presented in the article hold mutatis mutandis (this will be made more clear in Section 6). We impose the tacit condition on our generic classes $\mathcal{M}\subseteq\mathcal{M}^{0}$ that $F_{p}\in\mathcal{M}$ for each $F\in\mathcal{M}$ and $p\in(0,1)$. This assumption holds for common choices of $\mathcal{M}$, such as $\mathcal{M}=\mathcal{M}^{q}$, $q\in[1,\infty]$. [^1] A notable exception, however, is $\mathcal{M}\setminus\mathcal{M}^{\infty}$. Note that $F_{p}=G_{p}$ if and only if $F(x)=G(x)$ for all $x\geqslant\mathrm{VaR}_{p}^{+}(F)$.
@@
For $p\in(0,1)$, a risk measure $\rho\colon\mathcal{M}\to\mathbb{R}$ is a $p$-tail risk measure if $\rho(F)=\rho(G)$ for all $F,G\in\mathcal{M}$ satisfying $F_{p}=G_{p}$. A risk measure $\rho\colon\mathcal{M}\to\mathbb{R}$ is tail relevant if it is a $p$-tail risk measure for some $p\in(0,1)$. In this case, we simply call it a tail risk measure.
[…]
@@
and such a $\rho^{*}$ is unique on the set $\mathcal{M}^{*}$ of distributions in $\mathcal{M}$ whose support is bounded from below (Proposition 1 of Liu and Wang (2021)). Conversely, to obtain a $p$-tail risk measure, $p\in(0,1)$, it suffices to specify a generic risk measure $\rho^{*}$ and to define $\rho$ via (2.3). We will define $(\rho,\rho^{*})$ as a pair, and study their joint properties.
@@ 3 Elicitability, identifiability and convex level sets
    A functional $T\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k})$, is $\mathcal{M}$-conditionally elicitable with a functional $T^{\prime}\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k^{\prime}})$ if $T^{\prime}$ is $\mathcal{M}$-elicitable and $T$ is $\mathcal{M}(r,T^{\prime})$-elicitable for any $r\in\mathbb{R}^{k^{\prime}}$, where
    $\mathcal{M}(r,T^{\prime})=\{F\in\mathcal{M}:r\in T^{\prime}(F)\}.$
@@
The mean functional is $\mathcal{M}^{1}$-identifiable with the strict $\mathcal{M}^{1}$-identification function $V(x,y)=x-y$, $x,y\in\mathbb{R}$. Under mild conditions, any other strict $\mathcal{M}^{1}$-identification function is of the form $h(x)(x-y)$, where $h\neq 0$ (Dimitriadis et al., 2023). Moreover, under similar conditions, any strictly $\mathcal{M}^{1}$-consistent scoring function for the mean is given by
@@
[…]
@@
    A functional $T\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k})$ satisfies the CxLS property on $\mathcal{M}$ if for all $F_{0},F_{1}\in\mathcal{M}$ and for all $\lambda\in(0,1)$ such that $(1-\lambda)F_{0}+\lambda F_{1}\in\mathcal{M}$ it holds that
    $T(F_{0})\cap T(F_{1})\subseteq T((1-\lambda)F_{0}+\lambda F_{1}).$
@@
    A functional $T\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k})$ satisfies the CxLS* property on $\mathcal{M}$ if for all $F_{0},F_{1}\in\mathcal{M}$ and for all $\lambda\in(0,1)$ such that $(1-\lambda)F_{0}+\lambda F_{1}\in\mathcal{M}$ it holds that
    $T(F_{0})\cap T(F_{1})\neq\emptyset\implies T(F_{0})\cap T(F_{1})=T((1-\lambda)F_{0}+\lambda F_{1}).$
@@
Obviously, the CxLS* property implies the CxLS property. They both generalize the CxLS property for singleton-valued functionals, for which they coincide. Importantly for us, elicitability implies the CxLS* property (Fissler et al., 2021, Proposition 3.4), and similarly, identifiability implies the CxLS property (Proposition B.4 of the preprint version of Fissler and Hoga (2023)). This necessity has already partially been established by Osband (1985) and Gneiting (2011a). Under additional regularity conditions and for real-valued functionals, Steinwart et al. (2014) also showed the sufficiency of the CxLS property for elicitability and identifiability.
@@ 4 Identifiability results
$\mathcal{M}^{\rm c}_{p}$ =$=\left\{F\in\mathcal{M}^{0}:F^{-1}\text{ is continuous at }p\right\},\quad\text{ for }p\in(0,1);$
$\mathcal{M}_{\geqslant r}$ =$=\left\{F\in\mathcal{M}^{0}:F^{-1}(0)\geqslant r\right\},\quad\text{ for }r\in\mathbb{R};$
$\mathcal{M}_{(p)}$ =$=\left\{F\in\mathcal{M}^{0}:F(F^{-1}(p))=p\right\},\quad\text{ for }p\in[0,1).$
@@
    For any $r\in\mathbb{R}$ it holds that if $(Q_{p},\rho)$ is $\mathcal{M}_{(p)}\cap\big{(}\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}\big{)}$-identifiable with a strict identification function
    $\big{(}\mathds{1}_{\{y\leqslant v\}}-p,V(v,x,y)\big{)},~{}~{}~{}~{}v,x,y\in\mathbb{R},$
    for some function $V\colon\mathbb{R}^{3}\to\mathbb{R}$, then $\rho^{*}$ is $\mathcal{M}_{\geqslant r}\cap\mathcal{M}_{(0)}\cap\mathcal{M}$-identifiable with the strict identification function
@@
    Let $V$ be a strict $\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$-identification function for $\rho$ and $V_{r}^{*}$ as given in (4.1) for some $r\in\mathbb{R}$. Choose some $G\in\mathcal{M}_{\geqslant r}\cap\mathcal{M}$ and define
    $F(y)=\mathds{1}_{\{y\geqslant r\}}\big{(}(1-p)G(y)+p\big{)}=(1-p)G(y)+p\mathds{1}_{\{y\geqslant r\}}.$
    Due to the assumed identifiability of $\rho$ and since $F_{p}=G$, we get
    $\rho^{*}(G)=\rho^{*}(F_{p})=\rho(F)$ =$=\left\{x\in\mathbb{R}:\int V(x,y)\,\mathrm{d}F(y)=0\right\}$
    =$=\left\{x\in\mathbb{R}:\int V(x,y)\,\mathrm{d}\left(\mathds{1}_{\{y\geqslant r\}}(1-p)G(y)\right)+pV(x,r)=0\right\}$
    =$=\left\{x\in\mathbb{R}:\int\big{(}(1-p)V(x,y)+pV(x,r)\big{)}\,\mathrm{d}G(y)=0\right\}$
    =$=\left\{x\in\mathbb{R}:\int V^{*}_{r}(x,y)\,\mathrm{d}G(y)=0\right\},$
@@
    Let $\big{(}\mathds{1}_{\{y\leqslant v\}}-p,V(v,x,y)\big{)}$, $v,x,y\in\mathbb{R}$, be a strict $\mathcal{M}_{(p)}\cap\big{(}\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}\big{)}$-identification function for $(Q_{p},\rho)$ and $V_{r}^{*}$ given in (4.2) for some $r\in\mathbb{R}$. For any $F\in\mathcal{M}_{(p)}\cap\big{(}\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}\big{)}$ and $v\in Q_{p}(F)$ it holds that
    $\int V(v,x,y)\,\mathrm{d}F(y)=0\quad\text{ if and only if }\quad x=\rho(F).$
    Let $G\in\mathcal{M}_{\geqslant r}\cap\mathcal{M}_{(0)}\cap\mathcal{M}$. Then, as in part (i), define $F(y)=\mathds{1}_{\{y\geqslant r\}}\big{(}(1-p)G(y)+p\big{)}$, resulting in $F\in\mathcal{M}_{(p)}\cap\big{(}\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}\big{)}$ and $r=\mathrm{VaR}_{p}^{-}(F)\in Q_{p}(F)$. Again, $F_{p}=G$, which implies
    $\rho^{*}(G)=\rho^{*}(F_{p})=\rho(F)$ =$=\left\{x\in\mathbb{R}:\int V(r,x,y)\,\mathrm{d}F(y)=0\right\}$
    =$=\left\{x\in\mathbb{R}:\int V(r,x,y)\,\mathrm{d}\left(\mathds{1}_{\{x\geqslant r\}}(1-p)G(x)\right)+pV(r,x,r)=0\right\}$
    =$=\left\{x\in\mathbb{R}:\int\big{(}(1-p)V(r,x,y)+pV(r,x,r)\big{)}\,\mathrm{d}G(x)=0\right\},$
@@
We illustrate the part (i) of Proposition 4.1. Let $\mathcal{M}$ be the class of continuous distribution functions which are strictly increasing on their support. For any $\alpha\in(0,1)$, $p\in(0,\alpha]$ and $r\in\mathbb{R}$, $\mathrm{VaR}_{\alpha}^{-}$ is $\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$-identifiable with the strict identification function $V(x,y)=\mathds{1}_{\{y\leqslant x\}}-\alpha$. Moreover, $\mathrm{VaR}_{\alpha}^{-}$ is a $p$-tail risk measure with generator $\mathrm{VaR}_{\alpha^{*}}^{-}$, where $\alpha^{*}=(\alpha-p)/(1-p)$. Then, we (4.1) yields $V_{r}^{*}(x,y)=(1-p)\big{(}\mathds{1}_{\{y\leqslant x\}}-\alpha\big{)}+p\big{(}\mathds{1}_{\{y\leqslant r\}}-\alpha\big{)}$. Indeed, for any $F\in\mathcal{M}_{\geqslant r}\cap\mathcal{M}$ it holds that

### v3

@@ Abstract
Tail risk measures are fully determined by the distribution of the underlying loss beyond its quantile at a certain level, with Value-at-Risk, Expected Shortfall and Range Value-at-Risk being prime examples. They are induced by law-based risk measures, called their generators, evaluated on the tail distribution. This paper establishes joint identifiability and elicitability results of tail risk measures together with the corresponding quantile, provided that their generators are identifiable and elicitable, respectively. As an example, we establish the joint identifiability and elicitability of the tail expectile together with the quantile. The corresponding consistent scores constitute a novel class of weighted scores, nesting the known class of scores of Fissler and Ziegel for the Expected Shortfall together with the quantile. For statistical purposes, our results pave the way to easier model fitting for tail risk measures via regression and the generalized method of moments, but also model comparison and model validation in terms of established backtesting procedures.
@@ 1 Introduction
In view of the practical importance of tail risk measures, this paper studies their identifiability and elicitability. A risk measure is identifiable if it is the unique zero (root) of an expected identification function. Similarly, it is elicitable if it is the unique minimizer of an expected score; see Definitions 3.1 and 3.2. As such, identifiability and elicitability facilitate Z- and M-estimation (Huber and Ronchetti, 2009). More generally, model fitting can be pursued via the generalized method of moments (Newey and McFadden, 1994) or regression (Dimitriadis et al., 2024), also using recent tools from machine learning; see Fissler et al. (2023) and Cont et al. (2025) for examples. For model comparison and model validation, commonly subsumed under the umbrella term backtesting, identifiability and elicitability are crucial (Nolde and Ziegel, 2017; Fissler et al., 2016; Wang et al., 2025).
@@ 2.1 Notation
Let $\mathcal{M}^{0}$ be the set of all Borel probability distributions on $\mathbb{R}$ and let $\mathcal{M}^{q}\subseteq\mathcal{M}^{0}$, $q\in[1,\infty)$, be the subset of all distributions with a finite $q$th moment. Moreover, $\mathcal{M}^{\infty}$ denotes the set of all compactly supported distributions. For the ease of presentation, we identify distributions with the corresponding cumulative distribution functions; that is, $F(x)$ is understood as $F((-\infty,x])$ for $x\in\mathbb{R}$. For $F\in\mathcal{M}^{0}$, we define the left-continuous generalized inverse (left-quantile) as
@@
and in addition set $F^{-1}(0)=\inf\{x\in\mathbb{R}:F(x)>0\}$. For a function $F$ on a subset of $\mathbb{R}$, its left and right limits at $x$ are given by $F(x-)=\lim_{y\uparrow x}F(y)$ and $F(x+)=\lim_{y\downarrow x}F(y)$, assuming that the limits are well-defined. For $x\in\mathbb{R}$, denote by $\delta_{x}$ the point-mass probability measure at $x$. For simplicity, $\int$ means integration on $\mathbb{R}$ unless otherwise specified. Throughout this article, if we do not specify $\mathcal{M}$, then the statements hold for any set $\mathcal{M}\subseteq\mathcal{M}^{0}$ such that the functional at consideration is well-defined and finite on $\mathcal{M}$.
@@ 2.2 Risk measures
A risk measure is a mapping $\rho\colon\mathcal{M}\to$$\infty$$\infty$], where the set $\mathcal{M}\subseteq\mathcal{M}^{0}$ is the domain of $\rho$. The typical interpretation of a risk measure is that its argument, $F\in\mathcal{M}$, represents the distribution of a random loss from some financial position, and that $\rho(F)$ represents the amount of capital required to hold to make the financial position acceptable. In the literature, risk measures are often defined as mappings from a set of random variables, instead of $\mathcal{M}$, to the extended real line. When such risk measures are law-based, they one-to-one correspond to risk measures under our definition. We explain in Appendix A details on risk measures and their desirable properties.
The two most popular classes of risk measures used in banking and insurance practice are the Value-at-Risk (VaR) and the Expected Shortfall (ES). For a confidence level $p$, the right-quantile ($\mathrm{VaR}^{+}_{p}$) and the left-quantile ($\mathrm{VaR}_{p}^{-}$), are defined for $F\in\mathcal{M}^{0}$ as
$\mathrm{VaR}^{+}_{p}(F)$ $=\inf\{x\in\mathbb{R}:F(x)>p\}=F^{-1}(p+),~~~p\in[0,1);$
$\mathrm{VaR}^{-}_{p}(F)$ $=\inf\{x\in\mathbb{R}:F(x)\geqslant p\}=F^{-1}(p),~~~p\in(0,1].$
@@
As a compromise between the robustness of VaR and and the coherence of ES, Cont et al. (2010) introduced the so called Range Value-at-Risk (RVaR). For $0<p<q<1$, it is defined as
$\mathrm{RVaR}_{p,q}(F)=\frac{1}{q-p}\int_{p}^{q}\mathrm{VaR}^{+}_{r}(F)\,\mathrm{d}r\in\mathbb{R},\quad F\in\mathcal{M}^{0}.$ (2.2)
Clearly, for the boundary cases of $q=1$ or $p=q$, by taking limits, $\mathrm{RVaR}_{p,q}$ coincides with $\mathrm{ES}_{p}$ or $\mathrm{VaR}_{p}^{+}$, respectively.
@@ 2.3 Tail distribution and tail risk measures
Invoking our sign convention that a distribution $F\in\mathcal{M}^{0}$ is a loss distribution, it is clear that the right tail is the region of interest from a risk measurement and management perspective (McNeil et al., 2015). Clearly, with another sign convention and from a mathematical perspective, we could similarly consider the left tail of the distribution, and all the results presented in the article hold mutatis mutandis (this will be made more clear in Section 6). We impose the tacit condition on our generic classes $\mathcal{M}\subseteq\mathcal{M}^{0}$ that $F_{p}\in\mathcal{M}$ for each $F\in\mathcal{M}$ and $p\in(0,1)$. This assumption holds for common choices of $\mathcal{M}$, such as $\mathcal{M}=\mathcal{M}^{q}$, $q\in[1,\infty]$. [^1] A notable exception, however, is $\mathcal{M}\setminus\mathcal{M}^{\infty}$. Note that $F_{p}=G_{p}$ if and only if $F(x)=G(x)$ for all $x\geqslant\mathrm{VaR}_{p}^{+}(F)$.
@@
For $p\in(0,1)$, a risk measure $\rho\colon\mathcal{M}\to\mathbb{R}$ is a $p$-tail risk measure if $\rho(F)=\rho(G)$ for all $F,G\in\mathcal{M}$ satisfying $F_{p}=G_{p}$. A risk measure $\rho\colon\mathcal{M}\to\mathbb{R}$ is a tail risk measure if it is a $p$-tail risk measure for some $p\in(0,1)$.
It is immediate from Definition 2.1 that for $p,q\in(0,1)$ where $p<q$, $\mathrm{VaR}^{+}_{p}$, $\mathrm{ES}_{p}$ and $\mathrm{RVaR}_{p,q}$ are $r$-tail risk measures for $r\in(0,p]$, but not for $r\in(p,1)$; $\mathrm{VaR}^{-}_{p}$ is an $r$-tail risk measure only for $r\in(0,p)$. The expectation is not a tail risk measure. An important property of tail risk measures is that they can be generated by other risk measures. Precisely, for any $p\in(0,1)$ and $p$-tail risk measure $\rho$ on $\mathcal{M}$, there exists a risk measure $\rho^{*}$ on $\mathcal{M}$, called the $p$-generator of $\rho$, satisfying
@@
and such a $\rho^{*}$ is unique on the set $\mathcal{M}^{*}$ of distributions in $\mathcal{M}$ whose support is bounded from below (Proposition 1 of Liu and Wang, 2021). Conversely, to obtain a $p$-tail risk measure, $p\in(0,1)$, it suffices to specify a generic risk measure $\rho^{*}$ and to define $\rho$ via (2.3). We will define $(\rho,\rho^{*})$ as a pair, and study their joint properties.
@@
A slightly different approach of measuring tail risk is through the loss part of the random variables instead of the part above a quantile level; see Jarrow (2002) and Cont et al. (2013). We focus on the formulation in Definition 2.1 as it includes the most popular risk measures VaR and ES and many other examples of interest.
@@ 3 Elicitability, identifiability and convex level sets
    A functional $T\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k})$, is $\mathcal{M}$-conditionally elicitable with a functional $T^{\prime}\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k^{\prime}})$ if $T^{\prime}$ is $\mathcal{M}$-elicitable and $T$ is $\mathcal{M}(r,T^{\prime})$-elicitable for any $r\in\mathbb{R}^{k^{\prime}}$, where M(r,T’) = {F∈M: r∈T’(F)}.
@@
The mean functional is $\mathcal{M}^{1}$-identifiable with the strict $\mathcal{M}^{1}$-identification function V(x,y)=xyV(x,y)=x-y, $x,y\in\mathbb{R}$. Under mild conditions, any other strict $\mathcal{M}^{1}$-identification function is of the form h(x)(xy)h(x)(x-y), where $h\neq 0$ (Dimitriadis et al., 2024). Moreover, under similar conditions, any strictly $\mathcal{M}^{1}$-consistent scoring function for the mean is given by
@@
$\begin{aligned} S(v,x,y)&=\,\mathds{1}_{\{y>v\}}g(y)+\left(\mathds{1}_{\{y\leqslant v\}}-p\right)g(v)\\ &\qquad+\phi^{\prime}(x)\Big(x-\tfrac{1}{1-p}\big[\mathds{1}_{\{y>v\}}y+(\mathds{1}_{\{y\leqslant v\}}-p)v\big]\Big)-\phi(x)+a(y),\quad v,x,y\in\mathbb{R},\end{aligned}$ (3.4)
@@
Example 3.6.
Similarly to $\mathrm{ES}_{p}$, Fissler and Ziegel (2021b) showed that $\mathrm{RVaR}_{p,q}$, $0<p<q<1$, is not identifiable or elicitable on sufficiently rich classes $\mathcal{M}$. In the same vein as before, they established the $\mathcal{M}_{(p)}\cap\mathcal{M}_{(q)}$-identifiability of the triplet $(Q_{p},Q_{q},\mathrm{RVaR}_{p,q})$ via the identification function
$V(v_{1},v_{2},x,y)=\begin{pmatrix}\mathds{1}_{\{y\leqslant v_{1}\}}-p\\ \mathds{1}_{\{y\leqslant v_{2}\}}-q\\ x-\tfrac{1}{q-p}\Big[\mathds{1}_{\{v_{1}<y\leqslant v_{2}\}}y+(\mathds{1}_{\{y\leqslant v_{1}\}}-p)v_{1}-(\mathds{1}_{\{y\leqslant v_{2}\}}-q)v_{2}\Big]\end{pmatrix}$ (3.5)
for $v_{1},v_{2},x,y\in\mathbb{R}$. Moreover, strictly $\mathcal{M}^{0}$-consistent scoring functions for this triplet are of the form
$\begin{aligned} &S(v_{1},v_{2},x,y)\\ &=\,\mathds{1}_{\{y>v_{1}\}}g_{1}(y)+\left(\mathds{1}_{\{y\leqslant v_{1}\}}-p\right)g_{1}(v_{1})+\mathds{1}_{\{y>v_{2}\}}g_{2}(y)+\left(\mathds{1}_{\{y\leqslant v_{2}\}}-q\right)g_{2}(v_{2})\\ &\quad~+\phi^{\prime}(x)\Big(x-\tfrac{1}{q-p}\Big[\mathds{1}_{\{v_{1}<y\leqslant v_{2}\}}y+(\mathds{1}_{\{y\leqslant v_{1}\}}-p)v_{1}-(\mathds{1}_{\{y\leqslant v_{2}\}}-q)v_{2}\Big]\Big)-\phi(x)+a(y)\end{aligned}$ (3.6)
for $v_{1},v_{2},x,y\in\mathbb{R}$, where $\phi\colon\mathbb{R}\to\mathbb{R}$ is strictly convex with subgradient $\phi^{\prime}$, such that for all $x$ the functions $v_{1}\mapsto g_{1}(v_{1})-\frac{1}{q-p}\phi^{\prime}(x)v_{1}$ and $v_{2}\mapsto g_{2}(v_{2})+\frac{1}{q-p}\phi^{\prime}(x)v_{2}$ are strictly increasing, and where $a\colon\mathbb{R}\to\mathbb{R}$ is a function such that the expectation $\int S(v_{1},v_{2},x,y)\,\mathrm{d}F(y)$ is finite for all $F\in\mathcal{M}^{0}$.
@@
    A functional $T\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k})$ satisfies the CxLS property on $\mathcal{M}$ if for all $F_{0},F_{1}\in\mathcal{M}$ and for all $\lambda\in(0,1)$ such that $(1-\lambda)F_{0}+\lambda F_{1}\in\mathcal{M}$ it holds that T(F_0)∩T(F_1) ⊆T((1-λ)F_0+λF_1).
@@
    A functional $T\colon\mathcal{M}\to\mathcal{P}(\mathbb{R}^{k})$ satisfies the CxLS* property on $\mathcal{M}$ if for all $F_{0},F_{1}\in\mathcal{M}$ and for all $\lambda\in(0,1)$ such that $(1-\lambda)F_{0}+\lambda F_{1}\in\mathcal{M}$ it holds that T(F_0)∩T(F_1) ≠∅⟹T(F_0)∩T(F_1) = T((1-λ)F_0+λF_1).
@@
Obviously, the CxLS* property implies the CxLS property. They both generalize the CxLS property for singleton-valued functionals, for which they coincide. Importantly for us, elicitability implies the CxLS* property (Fissler et al., 2021, Proposition 3.4), and similarly, identifiability implies the CxLS property (Proposition B.4 of the preprint version of Fissler and Hoga, 2024). This necessity has already partially been established by Osband (1985) and Gneiting (2011a). Under additional regularity conditions and for real-valued functionals, Steinwart et al. (2014) also showed the sufficiency of the CxLS property for elicitability and identifiability.
@@ 4 Identifiability results
$\mathcal{M}^{\rm c}_{p}$ $=\left\{F\in\mathcal{M}^{0}:F^{-1}\text{ is continuous at }p\right\},\quad\text{ for }p\in(0,1);$
$\mathcal{M}_{\geqslant r}$ $=\left\{F\in\mathcal{M}^{0}:F^{-1}(0)\geqslant r\right\},\quad\text{ for }r\in\mathbb{R};$
$\mathcal{M}_{(p)}$ $=\left\{F\in\mathcal{M}^{0}:F(F^{-1}(p))=p\right\},\quad\text{ for }p\in[0,1).$
@@
[…]
@@
[…]
    $\rho^{*}(G)=\rho^{*}(F_{p})=\rho(F)$ $=\left\{x\in\mathbb{R}:\int V(x,y)\,\mathrm{d}F(y)=0\right\}$
    $=\left\{x\in\mathbb{R}:\int V(x,y)\,\mathrm{d}\left(\mathds{1}_{\{y\geqslant r\}}(1-p)G(y)\right)+pV(x,r)=0\right\}$
    $=\left\{x\in\mathbb{R}:\int\big((1-p)V(x,y)+pV(x,r)\big)\,\mathrm{d}G(y)=0\right\}$
    $=\left\{x\in\mathbb{R}:\int V^{*}_{r}(x,y)\,\mathrm{d}G(y)=0\right\},$
@@
    Let $\big(\mathds{1}_{\{y\leqslant v\}}-p,V(v,x,y)\big)$, $v,x,y\in\mathbb{R}$, be a strict $\mathcal{M}_{(p)}\cap\big(\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}\big)$-identification function for $(Q_{p},\rho)$ and $V_{r}^{*}$ given in (4.2) for some $r\in\mathbb{R}$. For any $F\in\mathcal{M}_{(p)}\cap\big(\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}\big)$ and $v\in Q_{p}(F)$ it holds that ∫V(v,x,y) dF(y)=0   if and only if  x=ρ(F). Let $G\in\mathcal{M}_{\geqslant r}\cap\mathcal{M}_{(0)}\cap\mathcal{M}$. Then, as in part (i), define $F(y)=\mathds{1}_{\{y\geqslant r\}}\big((1-p)G(y)+p\big)$, resulting in $F\in\mathcal{M}_{(p)}\cap\big(\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}\big)$ and $r=\mathrm{VaR}_{p}^{-}(F)\in Q_{p}(F)$. Again, $F_{p}=G$, which implies
    $\rho^{*}(G)=\rho^{*}(F_{p})=\rho(F)$ $=\left\{x\in\mathbb{R}:\int V(r,x,y)\,\mathrm{d}F(y)=0\right\}$
    $=\left\{x\in\mathbb{R}:\int V(r,x,y)\,\mathrm{d}\left(\mathds{1}_{\{x\geqslant r\}}(1-p)G(x)\right)+pV(r,x,r)=0\right\}$
    $=\left\{x\in\mathbb{R}:\int\big((1-p)V(r,x,y)+pV(r,x,r)\big)\,\mathrm{d}G(x)=0\right\},$
@@
We illustrate the part (i) of Proposition 4.1. Let $\mathcal{M}$ be the class of continuous distribution functions which are strictly increasing on their support. For any $\alpha\in(0,1)$, $p\in(0,\alpha]$ and $r\in\mathbb{R}$, $\mathrm{VaR}_{\alpha}^{-}$ is $\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$-identifiable with the strict identification function $V(x,y)=\mathds{1}_{\{y\leqslant x\}}-\alpha$. Moreover, $\mathrm{VaR}_{\alpha}^{-}$ is a $p$-tail risk measure with generator $\mathrm{VaR}_{\alpha^{*}}^{-}$, where $\alpha^{*}=(\alpha-p)/(1-p)$. Then, (4.1) yields $V_{r}^{*}(x,y)=(1-p)\big(\mathds{1}_{\{y\leqslant x\}}-\alpha\big)+p\big(\mathds{1}_{\{y\leqslant r\}}-\alpha\big)$. Indeed, for any $F\in\mathcal{M}_{\geqslant r}\cap\mathcal{M}$ it holds that